# Chaos-based reinforcement learning with TD3

### v1

@@ 4.7 Exploration with random number layer
The above verification confirmed that learning proceeds to ignore the reservoir’s output if the reservoir is too chaotic. To verify this result further, we investigated learning when a random vector of independent and identically distributed random numbers replaces the reservoir. Specifically, we replaced the reservoir with a uniform random vector sampled from $[-1,1]$ and conducted learning under this condition. In this setting, the random vector serves as the exploration component but is worthless as information about the agent’s state. Figure 14 shows the results of this experiment. This figure shows that the agent succeeded in learning the task even when a random number vector replaces the reservoir, although the number of steps required for learning tends to increase. The learning results for the readout weights under this situation show that the bypass weights are significantly larger than the weights from the random layer. This result seems to be because the random layer’s output is worthless for accomplishing the task and suggests that a similar phenomenon occurs when the spectral radius of the reservoir is too high.
@@
Fig. 14: Result when a random vector is used instead of a reservoir. The definitions of line colors are the same as in Fig. 5, 6, and 11.
@@
Fig. 15: Result of the goal change task when a random vector is used instead of a reservoir. The definitions of the line colors are the same as in Fig. 5, 6, and 11.
5 Conclusion
This study introduced TD3 as a learning algorithm for chaos-based reinforcement learning (CBRL) and revealed several characteristics.
@@ 5 Conclusion
It was confirmed that the TD3-CBRL agent can learn the goal task. This suggests that TD3 is a good candidate as a learning algorithm for CBRL, which had not been well-established in previous studies. Although a regular TD3 agent succeeded in learning the goal task, the agent failed to learn it without external exploration noise, despite the simplicity of the task. On the other hand, the TD3-CBRL agent, whose actor model has chaotic dynamics, can learn without exploration with random noises. This comparison indicates that the internal chaotic dynamics of the reservoir contributes to the agent’s exploratory behavior.
We also observed the agent’s adaptation to the environmental change of altering the goal position during the learning process and found that the agent resumed chaotic exploration and successfully re-learned. Furthermore, our results indicate that there is an appropriate range for the strength of internal chaos for the TD3-CBRL agent to have the flexibility to autonomously switch between exploration and exploitation modes. In particular, it was found that when the spectral radius is too large, the weights from the bypass become excessively large. Then, the inflated weights negatively affect the flexibility of switching between exploration and exploitation. These results indicate that the appropriate choice of spectral radius is essential in designing reservoir networks in CBRL to achieve a proper balance between exploration and exploitation.
We tested the TD3-CBRL agent when the reservoir network, the source of the exploration dynamics, was replaced by a random vector. The results showed that the agent could learn the goal task, but the bypass weights became significantly large, and the agent failed to learn the goal change task. These results indicate that the reservoir, which generates the dynamics in CBRL, does not act just as a mere noise source. While the random vectors do not reflect any external input, the reservoir creates a variety of background activities in the system incorporating the input’s influence. This activity seems to provide the variability necessary for exploration and to be a resource that drives new exploratory behavior when encountering unknown experiences.
@@
Various studies have shown that neural populations operate in a critical state [61, 62, 63, 64]. Reservoir network performance has also been shown to be optimized at the edges of chaos [57, 58]. In addition, a study in which ESNs performed chaos-based exploration learning also showed results suggesting that exploration and exploitation are balanced at the edges of chaos [43]. In this study, the spectral radius was not optimal around $g=1$, where the reservoir dynamics is typically on the edge of chaos. One hypothesis for explaining this result is that the spectral radius of the subsystem, the reservoir, needed to be larger to place the entire system’s dynamics, including the environment, at the edge of chaos. It is important to verify this by focusing on the entire system’s behavior, including the interaction between the CBRL agent and the environment.
This study introduced TD3 to CBRL and validated its ability with a simple task to investigate learning and re-learning in detail. In previous studies and this one, CBRL has yet to learn difficult benchmark tasks used for deep reinforcement learning research, such as Atari game tasks and nonlinear control tasks using MUJOCO. In future works, improving the model to learn more difficult tasks is necessary. A wider selection of tasks will allow us to examine CBRL from newer perspectives.
@@
Introducing the new reservoir structure proposed to extend its performance is useful to improve the learning performance of CBRL. Reservoirs do not perform well with high-dimensional input such as images. Several studies have proposed methods that use untrained CNNs for feature extraction [65, 66]. Introducing these methods can enable CBRL agents to learn tasks with high-dimensional input, such as raw images. Structural improvement that constructs reservoirs in multiple layers [67, 68, 69] and methods that reduce the model size by using multi-step reservoir outputs as input to readouts [70] have been proposed. It is worth verifying the use of these new reservoir techniques to improve the performance of CBRL agents.
@@
Improvements in learning algorithms are also worth considering. Sensitivity adjustment learning (SAL), which adjusts the chaoticity of neurons based on ”sensitivity,” has been proposed [71]. This method can modify the recurrent weights while keeping the chaoticity of the RNN. Using SAL to maintain chaoticity and learning with Backpropagation Through Time may allow CBRL agents to learn more difficult tasks. Self-modulated reservoir computing (SM-RC) that extends the reservoir network’s ability by dynamically changing the characteristics of the reservoir and attention to the input through a self-modulated function has been proposed [72]. Since SM-RC can adjust its spectral radius, it is also expected that CBRL agents with SM-RC can learn to change their chaoticity and switch between exploration and exploitation states more dynamically.

### v2

@@ 4.6 Exploration with random number layer
The above verifications confirmed that learning proceeds to ignore the reservoir’s output and the experiences in the replay buffer stored before goal change negatively affect the re-learning ability if the reservoir is excessively chaotic. To verify this result further, we investigated learning when a random vector of independent and identically distributed random numbers replaces the reservoir. Specifically, we replaced the reservoir with a uniform random vector sampled from [1,1][-1,1] and conducted learning under this condition. This setup is designed to reproduce the characteristics of a reservoir with extremely strong chaoticity. In this setup, the random vector serves as the exploration component but is worthless as information about the agent’s state. Figure 13 shows the results of this experiment. This figure shows that the agent succeeded in learning the task even when a random number vector replaces the reservoir, although the number of steps required for learning tends to increase. The agent’s trajectory in the environment is noisy due to the influence of a random vector. The learning results for the readout weights under this situation show that the bypass weights are significantly larger than the weights from the random layer. This result seems to be because the random layer’s output is worthless as an input for accomplishing the task and suggests that a similar phenomenon occurs when the spectral radius of the reservoir is too high.
@@
Figure 15 shows the results when the replay buffer size is set to 64. This figure indicates that while the agent’s trajectory fluctuates significantly and the steps to the goal are numerous, it successfully learns to reach the new goal after the goal change. This result indicates that similar to the case with the reservoir, reducing the buffer size enables re-learning even when the model is highly irregular. These results suggest that storing meaningless vector information about the environment in the replay buffer hinders learning after environmental changes and slows down re-learning when $g$ of the reservoir is excessively large.
We observed that even when the reservoir exhibits strongly chaotic behavior, or the hidden layer is a random vector, the agent successfully re-learn if the replay buffer size is small and experiences before environmental changes are quickly removed. However, when the system is highly irregular, the trajectory toward the goal fluctuates significantly, negatively impacting performance. Additionally, experience replay is a crucial technique for improving sample efficiency, then it is undesirable to use a small buffer size. Therefore, appropriately setting the chaos level of the reservoir is critical to achieving both effective re-learning and improved learning performance in CBRL.
@@
Fig. 13: Result when a random vector is used instead of a reservoir. The definitions of line colors are the same as in Figs. 5, 6, and $1$.
@@
Fig. 14: Result of the goal change task when a random vector is used instead of a reservoir. The definitions of the line colors are the same as in Figs. 5, 6, and $1$.
(a) Learning curve.
(b) Readout weights.
(c) Agent trajectories.
Fig. 15: Result of the goal change task when a random vector is used instead of a reservoir and replay buffer size is set to 64. The definitions of the line colors are the same as in Figs. 5, 6, and $1$.
4.7 reservoir dynamics
The experimental results in Section 4.5 and 4.6 suggest that the output of the reservoir, which is stored in the replay buffer before the goal change, affects the capability of re-learning. In other words, it is considered that some reservoir characteristics other than the role as a source of exploration components contribute effectively to re-learning.
To investigate the role of the reservoir other than as a source of exploration components, we trained a CBRL agent by adding external random noise to the action output and examined the relationship between the spectral radius and learning performance. Figure 16(a) shows the results of this experiment. The figure shows that even without chaotic properties in the reservoir, the agent succeeded in learning the task when external exploration noise was used. Figure 16(b) shows the results of a similar verification for the goal change task. From the figure, we can confirm that there is an appropriate range for the spectral radius in this case. Interestingly, in cases where the task environment changes, learning performance decreases in the range where $g$ is small, even when external exploration noises are added to the actions. This result demonstrates that properly tuned chaotic reservoirs possess properties not found in random noise, which contribute to their effectiveness for re-learning.
(a) Goal task.
(b) Goal change task.
Fig. 16: Learning performance of TD3-CBRL agent with the external random noise added to the action outputs, varying spectral radius. The definitions of the line colors are the same as in Fig. 10. (a) shows the learning results for the goal task. (b) shows the learning results of the goal change task.
Figure 17(a) shows the trajectories of an agent with $g=0.9$ during the test phase of the goal change task, while Figure 17(b) shows the trajectories of the same agent when random noise is added to its actions during testing. From these figures, we observe that even when random noise is added to the action output, the agent continues to move towards the initial goal after the goal change, indicating that it has not learned to move towards the new goal.
(a) TD3-CBRL agent $(g=0.9)$ with external exploration random noise
(b) TD3-CBRL agent $(g=0.9)$ with external exploration random noise (with the noise in the test phase)
Fig. 17: Agent trajectories during the test of the goal change task. (a) shows the result of the TD3-CBRL agent with spectral radius $g$ of 0.9. (b) shows the behavior of the same agent as in (a) when the exploration by random noises is not stopped during the test. Each graph in the subfigures shows the test results conducted every 5000 training steps. The definitions of the line colors are the same as in Fig. 6.
Figure 18 shows the outputs of 5 reservoir neurons when the agent starts from $(10,18)$ during testing. From these figures, we can see that the reservoir state exhibits gradual dynamics when $g=0.9$. Notably, the output converges to a constant value after the initial transient and no longer changes. Even when random noise is added to the action output, we observe a similar trend with only slight perturbations in the output values. On the other hand, it fluctuates wildly when $g=2.2$.
(a) TD3-CBRL agent $(g=0.9)$ with external exploration random noise
(b) TD3-CBRL agent $(g=0.9)$ with external exploration noise even in the test phase
(c) TD3-CBRL agent $(g=2.2)$ (The same agent as in Fig. 6(a))
Fig. 18: Reservoir states when the agent starts from $(10,18)$ during testing. (a) shows the result with $g=0.9$. (b) shows the result of the same agent as (a) with external random noise during testing. (c) shows the result with $g=2.2$ (the same condition as in Section 4.4)
(a) $g=0.9$
(b) $g=0.9$ with external noise
(c) $g=2.2$
Fig. 19: PCA of reservoir states when the agent starts from $(10,18)$ during the test. A black star indicates the initial reservoir state. Blue, orange, and green trajectories show the reservoir state transitions in the test phase at the $20000$ steps, $20001$ steps (immediately after the goal change), and $50000$ steps (final step). The colored stars indicate the reservoir state at the end of the episode when the agent reaches the goal or the 200-step elapses. (a) shows the result of the TD3-CBRL agent $(g=0.9)$ with external exploration random noise. (b) shows the result of TD3-CBRL agent $(g=0.9)$ with external exploration random noise even in the test phase. (c) shows the result of TD3-CBRL agent $(g=2.2)$ (The same agent as in Fig. 6(a)).
To analyze these dynamics, we recorded the reservoir states obtained at all test phases during these learning processes of the goal change task and performed principal component analysis (PCA). Figure 19 shows the results of the dimensionality reduction of the reservoir states. The data in this figure is based on the reservoir states when the agent starts at $(10,18)$ and represents the test results after learning for $20000$ steps, $20001$ steps, and $50000$ steps. From the orange trajectory, we can observe how the reservoir state evolved after reaching the initial goal. Figure 19(a) shows that the agent with $g=0.9$ converges to a fixed-point after passing through the previous goal, and it can be seen that the trajectory converges to a slightly shifted point even after $50000$ steps. Figure 19(b) shows that even when random noise is added to the action output, the reservoir state does not escape the constructed attractor and fluctuates near the same attractor. These results indicate that the behavior of reaching the first learned goal constitutes such an attractor, which exerts an attractive force that cannot be escaped by random noise. On the other hand, Figure 19(c) shows that the reservoir’s chaoticity causes it to transition to a trajectory with irregular states after reaching the endpoint of the previously learned trajectory. The result at the 50000 steps shows that it has successfully constructed a trajectory towards the new goal. This result suggests that the dynamics of spontaneous activity due to chaotic state transitions, rather than convergence due to learned behavior, play an important role in autonomously shifting to new exploration. It is also noteworthy that the trajectories in Fig. 19 are close to each other in the initial steps. This suggests that the reservoir’s dynamics retain memory of the sequence of states observed from the environment.
Figure 20 presents the PCA results of the reservoir state for a TD3-CBRL agent with a reservoir of $g=2.2$, starting from two initial positions $(10,2)$ and $(10,18)$, during tests at $20000$ and $50000$ steps. This figure shows that trajectories from different start points follow distinct paths, while trajectories starting from the same point are close to each other in the initial steps. This indicates that the reservoir retains the memory of the state sequence from the starting position. Furthermore, the reservoir state trajectory at the $50000$ steps transits from a point close to the point when the agent reached the initial goal before the goal change to a new trajectory that heads towards the new goal while maintaining temporal continuity. This suggests that the reservoir’s short-term memory ability ensures the continuity of experiences before and after the environmental change stored in the replay buffer, which in turn positively affects the re-learning performance. Such spatio-temporally correlated state trajectories do not appear when using random number vectors. It can be considered that the absence of the correlation causes the difference in performance between the chaotic reservoir and the random vector in the re-learning task.
Fig. 20: PCA of reservoir states of TD3-CBRL agent $(g=2.2)$ (The same agent as in Fig. 6(a)) during the test. A black star indicates the initial reservoir state. Blue and orange trajectories show the results at the $20000$ steps and $50000$ steps starting from $(10,2)$. Green and red trajectories show the results at the $20000$ steps and $50000$ steps starting from $(10,18)$.
4.8 Learning performance and limitation
Although there are some variations in the definitions of actions and states, previous CBRL research has typically employed simple goal tasks similar to the one used in this paper. This is because previous CBRL algorithms had limitations in learning more challenging tasks. To investigate the learning performance of a CBRL agent with TD3 on relatively more difficult tasks, we conducted experiments using MuJoCo continuous control tasks [52] through OpenAI Gym [53], which are widely used as benchmarks in reinforcement learning. In this experiment, the critic network and actor network (the readout of the chaotic reservoir) were implemented as MLPs with two hidden layers, each comprising 128 ReLU activation nodes. The output layer’s neuron of the critic network was a linear activation node, while the neurons of the actor network were $\tanh$ nodes. The learning rate for both networks was set to $0.001$. A discount factor $\gamma$ was set to $0.99$, and the target network update time constant $\tau$ was set to $0.005$.
The learning results are shown in Figure 21. The figure shows that the agent exhibits comparable learning performance to the regular TD3 in learning tasks such as Pendulum, Inverted Pendulum, and Inverted Double Pendulum. On the other hand, the agent failed to learn more complex tasks, such as HalfCheetah, Walker2d, and Hopper. These results demonstrate that TD3-CBRL has successfully learned tasks that are more challenging than those in previous studies while simultaneously highlighting the current limitations of CBRL agents. Moreover, the figure reveals that setting the replay buffer size to 64 results in learning failure for all tasks. This emphasizes the importance of incorporating experience replay techniques for improving learning performance in CBRL.
(a) Pendulum
(b) Inverted Pendulum
(c) Inverted Double Pendulum
(d) Half Cheetah
(e) Walker2d
(f) Hopper
Fig. 21: Learning performance for the MuJoCo continuous control tasks of Open AI Gym. The vertical axis shows the average total reward from the results with ten different random number seeds. The horizontal axis shows the training steps. The blue, orange, and green lines show the average total rewards of TD3-CBRL, TD3-CBRL with a replay buffer size of 64, and regular TD3, respectively. Each shaded area shows its standard deviations. The results of regular TD3 were obtained from https://github.com/sfujim/TD3.
4.9 Partially observable Markov decision process (POMDP) task
Environments whose dynamics does not follow the Markov property and where the learning agent can observe only partial and incomplete information regarding the current state are formally described as Partially Observable Markov Decision Processes (POMDPs). Algorithms using RNNs have been shown to be one of the effective approaches for reinforcement learning in such environments [54, 55]. Since the CBRL agent uses an ESN as a source of chaotic dynamics, it possesses the potential to handle time-series processing. This section investigates whether the CBRL agent can use the short-term memory of the ESN to learn a POMDP task. The experiments in the previous sections have focused on the characteristics of the chaotic ESN as a source of exploration, using environments satisfying the Markov property. Here, we conduct learning on the “flickering goal task” to investigate whether the TD3-CBRL agent can learn POMDP tasks. The flickering problem is usually used to introduce partial observability to an environment in reinforcement learning tasks[56, 55]. Although the basic environmental rules in this task are the same as in the goal task, the agent can only observe the state of the environment with a probability of $p_{\rm{obs}}=0.5$ in this task, except in the first step of an episode. In other words, the input $\bm{u}_{t}$ is replaced by a zero vector with probability $(1-p_{\rm{obs}})$. To enable the critic network $Q$ to estimate state-action values from the observation sequence, the output of the ESN was concatenated with the inputs $\bm{u}_{t}$ to critic network $Q$. Furthermore, we adjusted the spectral radius to $g=1.2$ to ensure that the ESN fully exhibited not only the exploration capability but also the time-series processing capability and increased the number of learning steps to 50,000 to accommodate the increased task difficulty.
Figure 22(a) shows the learning curve for this experiment. This figure indicates that the number of steps to reach the goal decreases as learning progresses and that the TD3-CBRL agent is successfully learning the POMDP task. On the other hand, Fig. 22(b) shows the results when the ESN was not connected to the critic network Q. In this case, the agent fails to learn. This result demonstrates that time-series processing by the ESN is essential for predicting state-action values in the flickering goal task. Furthermore, as a comparison, Fig. 22(c) shows the learning results using a standard TD3 algorithm with added external exploration noise. This figure shows that regular TD3 agent fails to learn the flickering goal task, indicating that models without time-series processing capabilities cannot learn this task. These results suggest that TD3-CBRL agents have the potential to learn in situations where the current state provides insufficient information for decision-making.
(a) TD3-CBRL ($g=1.2$ and ESN is connected to $Q$)
(b) TD-CBRL ($g=1.2$)
(c) Regular TD3
Fig. 22: Learning curves of the flickering goal task. The definition of the line color is the same as in Fig. 5. (a) shows the learning results with TD3-CBRL ($g=1.2$) in which the ESN was connected to the critic network $Q$. (b) shows the learning results with TD3-CBRL ($g=1.2$). (c) shows the learning results with regular TD3.
In time-series processing with ESNs, satisfying the Echo State Property (ESP) is a necessary condition for stable and reliable operation [57]. The ESP depends on the spectral radius $g$ of the recurrent weight matrix of the ESN. We investigated the influence of $g$ varying from 0 to 2.0 in increments of 0.2. The results of this experiment are presented in Figure 23. The figure reveals a distinct peak in learning success probability centered around $g=1.0$. This result suggests that too small $g$ probably diminishes the ESN’s capacity to retain sufficient information from the observation history, limiting its memory of past input. In contrast, it suggests that too large $g$ induces highly sensitive temporal dynamics within the ESN, hindering the memory capability of past information necessary to learn the POMDP tasks.
These results suggest that while the TD3-CBRL agent with an ESN in its architecture can exhibit time-series processing capabilities, the trade-off between the short-term memory retention and the emergence of chaotic dynamics constrains the optimal range of $g$ to a relatively narrow one in learning POMDP tasks.
Fig. 23: Learning performance of the flickering goal task with varying the spectral radius of the ESN in the CBRL-TD3 agent. The definitions of the line colors are the same as in Fig. 10.
5 Discussion
@@
This study introduced TD3 as a learning algorithm for chaos-based reinforcement learning (CBRL) and revealed several characteristics. It was confirmed that the TD3-CBRL agent can learn the goal task. This suggests that TD3 is a good candidate as a learning algorithm for CBRL, which had not been well-established in previous studies. Although a regular TD3 agent succeeded in learning the goal task, the agent failed to learn it without external exploration noise, despite the simplicity of the task. On the other hand, the TD3-CBRL agent, whose actor model has chaotic dynamics, can learn without exploration with random noises. This comparison indicates that the internal chaotic dynamics of the reservoir contributes to the agent’s exploratory behavior.
As learning progressed, the agent was able to autonomously suppress its exploratory behavior and transition to an exploitation mode. Furthermore, we confirmed that the agent resumed chaotic exploration and successfully re-learned when faced with an environmental change, such as changing the goal location during the learning process. However, the regular TD3, where the model does not have a reservoir and exploration is driven by external random numbers, failed this re-learning task. This suggests that exploration based on the reservoir’s chaotic dynamics has properties that external random noise added to the action output does not possess.
The varying experiments revealed a suitable range of chaos strength in the agent’s model for the TD3-CBRL agent to switch between exploration and exploitation modes autonomously. Specifically, it was found that when the spectral radius is too large, or the reservoir states are almost completely random, the number of steps for re-learning becomes excessively large. This increase in required steps is mitigated by reducing the buffer size, although the learning performance tends to decrease. This suggests that the reservoir states stored in the replay buffer before the environmental change are affecting the re-learning process. In regular TD3, experiences stored before the rule change may be inappropriate for re-learning because they do not reflect the environment after the rule change. However, in CBRL, where the reservoir states are stored in the experience buffer, the reservoir states with an appropriately tuned $g$ can retain short-term memory of the state sequence from the start position. This enables ensuring the temporal continuity of experience before and after environmental changes. On the other hand, when the reservoir’s chaoticity is too strong or its output is almost completely random, it only provides irregular noise. It seems to be the reason why TD3-CBRL agents with excessive reservoir chaoticity fail to adapt to environmental changes and show reduced re-learning performance.
It has also been confirmed that the dynamics of the reservoir should not be convergent. In experiments where external noise was added to the TD3-CBRL for exploration, it was observed that even with the promotion of exploration by external noise, the agent failed to re-learn when $g$ was small. This result suggests that it is not enough for the reservoir to simply act as a state sequence memory layer. We observed the reservoir states in the re-learning task by applying PCA that reduces its dimensionality. The CBRL agent with convergent dynamics ($g=0.9$) failed to re-learn because its reservoir state could not escape from the fixed-point attractor. On the other hand, the agent with $g=2.2$ acquired a new behavior to reach the new goal after the reservoir exhibited a new irregular trajectory. This result suggests that chaotic reservoirs are essential for re-learning, as they generate rich dynamics that create a variety of background activities within the system, incorporating the influence of the input and enabling the formation of new attractors leading to the desired goal.
These results are consistent with the criticality hypothesis of the brain in neuroscience [58, 59]. The result observed in this study, that learning of the CBRL agents is optimized within a specific range of the spectral radius $g$, supports the functional advantages of this criticality in the context of reinforcement learning. While further investigation on task dependency is necessary, the chaos-based exploration exhibited more effective exploration capabilities than the external random number-based exploration in response to environmental changes in the goal task of this study. The agent’s internal chaotic dynamics can be regarded as a computational model of the brain’s spontaneous activity, which generates diverse patterns of activity. This activity seems to provide the fluctuations necessary for exploration, maintaining the temporal correlations of inputs in reservoir states, and being a resource that drives new exploratory behavior when encountering unknown experiences.
6 Conclusion and Future Work
CBRL is studied with the expectation of understanding a possible role of chaotic fluctuations on biological brains, realizing intelligent systems based on transient dynamics, and developing reinforcement learning algorithms that can optimize exploratory behavior through learning. This study extended the foundation of CBRL’s learning algorithm by incorporating TD3, and we confirmed the existence of an optimal range of chaos intensity for the best balance between exploration and exploitation. Furthermore, we demonstrated that our approach enables the learning of challenging tasks that were previously unmanageable, contributing to advancements in exploratory learning models utilizing chaotic fluctuations. However, further verification and overcoming challenges are necessary for the practical application of CBRL.
@@ 6 Conclusion and Future Work
Various studies have shown that neural populations operate in a critical state [58, 59, 60, 61]. Reservoir network performance has also been shown to be optimized at the edges of chaos [48, 49]. In addition, a study in which ESNs performed chaos-based exploration learning also showed results suggesting that exploration and exploitation are balanced at the edges of chaos [34]. It is a very important and promising future research direction to investigate how to quantify the chaoticity of CBRL agents. In this study, the spectral radius was not optimal around $g=1$, where the reservoir dynamics is typically on the edge of chaos. One hypothesis for explaining this result is that the spectral radius of the subsystem, the reservoir, needed to be larger to place the entire system’s dynamics, including the environment, at the edge of chaos. It is important to verify this by focusing on the entire system’s behavior, including the interaction between the CBRL agent and the environment. Furthermore, clarifying the differences between chaos-driven and random-noise exploration is a crucial challenge for evaluating the effectiveness of chaos-based exploration methods. To this end, investigating the distribution and entropy of behaviors and states and comparing them with the results obtained by conventional exploration using random noise is expected to provide valuable insights.
Experimental results of MuJoCo continuous control tasks indicate that TD3-CBRL can solve tasks that were difficult for the previous CBRL approaches. However, it was also found that TD3-CBRL is limited in its ability to learn tasks with high-dimensional state and action spaces. It is an important future problem to gain deeper insights by using various tasks. For example, tasks that require handling intermediate outputs could be considered. The optimal action in this study’s goal tasks is maximizing the output to reach the goal, which did not necessitate intermediate outputs, thus failing to fully utilize the advantages of TD3 capable of handling continuous action outputs. As tasks that require intermediate outputs, we may consider tasks that treat the agent’s acceleration as the action or tasks that impose penalties on large movements. Furthermore, it is necessary to conduct further investigations across a broader range of environments, rather than using a simple goal task, to evaluate how learning is affected by various conditions, such as tasks with a long time horizon, sparse rewards, multi-step planning, or the presence of observation noise. By improving the model to learn more complex tasks and conducting experiments with a wider range of tasks, we can examine the learning system based on exploration by internal chaotic fluctuations and interactions with the environment from a new perspective.
@@
Introducing the new reservoir structure proposed to extend its performance is useful to improve the learning performance of CBRL. Reservoirs do not perform well with high-dimensional input such as images. Several studies have proposed methods that use untrained convolutional neural networks for feature extraction [62, 63]. Introducing these methods can enable CBRL agents to learn tasks with high-dimensional input, such as raw images. Structural improvement that constructs reservoirs in multiple layers [64, 65, 66] and methods that reduce the model size by using multi-step reservoir states as input to readouts [67] have been proposed. It is worth verifying the use of these new reservoir techniques to improve the performance of CBRL agents.
@@
Improvements in learning algorithms are also worth considering. Sensitivity adjustment learning (SAL), which adjusts the chaoticity of neurons based on ”sensitivity,” has been proposed [68]. This method can modify the recurrent weights while keeping the chaoticity of the recurrent neural networks. Using SAL to maintain chaoticity and learning with Backpropagation Through Time may allow CBRL agents to learn more difficult tasks. Self-modulated reservoir computing (SM-RC) that extends the reservoir network’s ability by dynamically changing the characteristics of the reservoir and attention to the input through a self-modulated function has been proposed [69]. Since SM-RC can adjust its spectral radius, it is also expected that CBRL agents with SM-RC can learn to change their chaoticity and switch between exploration and exploitation states more dynamically. Furthermore, it is worth exploring approaches that combine the exploration driven by chaotic fluctuations with other algorithms aiming at efficient exploration, such as intrinsically motivated reinforcement learning and meta-learning, to guide the exploration in a more efficient direction.
Improving adaptability to non-stationary environments is a critical challenge in reinforcement learning, and a variety of methods have been proposed to address this issue. For instance, Steinparz et al. used intrinsic rewards to encourage exploration in response to environmental changes [70]. Canonaco et al. proposed NSD-RL that actively detects environmental changes to re-adapt [71], and Zhu et al. proposed a method that autonomously rebalances exploration and exploitation based on the model uncertainty of the Q-network [72]. While a detailed comparative analysis between these existing methods and CBRL remains a subject for future work, they are not mutually exclusive; combining them could yield synergistic effects. The inherent sensitivity of the chaotic dynamics utilized in CBRL may amplify the novelty of unknown states or facilitate the detection of subtle state changes arising from environmental shifts. Verifying the integration of CBRL with other algorithms designed for non-stationary environments is worthwhile future work for more efficient re-exploration and re-learning.